# Superconductivity from spin-canting fluctuations in rhombohedral graphene

### v2

@@ VI Discussion
    The phase diagram area populated by superconductivity shrinks upon increasing spin-orbit coupling. In our scenario the pairing glue—soft magnons—relies on spin-canting order in the normal state. As $\lambda_{I}$ increases, however, canting order eventually loses in favor of an Ising spin-valley-locked phase; recall Sec. II.1. It is thus natural to anticipate that the spin-canting phase occupies a progressively smaller region of the phase diagram as $\lambda_{I}$ rises, as also seen in Hartree-Fock studies [50, 51]. Our theory then predicts that superconductivity should also occupy a diminished area.
@@
    The critical temperature $T_{c}$ is enhanced upon increasing the spin-orbit coupling. This prediction follows directly from our analysis of the pairing interaction in Eq. (50), which contains a $\cos\theta_{0}\propto\lambda_{I}$ prefactor. This factor vanishes when $\lambda_{I}\rightarrow 0$, i.e., the divergence that drives two-magnon-mediated pairing in our scenario disappears. It is then natural that turning on $\lambda_{I}$ increases the optimal $T_{c}$ emerging from a spin-canted normal state, even though the latter’s phase space diminishes.
@@
Remarkably, Ref. 39 observed both of these trends in bilayer devices, where the strength of Ising spin-orbit interaction was controlled through engineering the twist angle at the WSe2-graphene interface [73, 74, 75]. Direct detection of the spin-canting order central to our theory (e.g., by SQUID measurements) poses arguably the simplest falsifiable check for our proposed explanation.
@@
Another nontrivial prediction concerns the dependence of superconductivity on carrier density. We showed that our proposed two-magnon exchange mechanism activates two $s$-wave pairing channels—one for the majority Fermi surfaces, and the other for the minority Fermi surfaces. Importantly, we find that the pairing interaction scales inversely proportional to the corresponding Fermi momenta, implying that pairing predominantly occurs in the minority species with a strength that—all other factors, e.g. density of states, being equal—grows when the minority carrier density decreases. This prediction indeed matches the generic behavior observed in bilayer and trilayer graphene where prominent superconducting phases appear in regions hosting small pockets [37, 38, 39, 40, 41, 42]; perhaps even more strikingly, in trilayers these regions abruptly terminate near Lifshitz transitions marking the disappearance of minority carriers [41, 42].
@@
Our pairing scenario also has ramifications for depairing by in-plane magnetic fields $B_{\parallel}$. Conventional in-plane-field depairing mechanisms include orbital effects and warping of the Fermi surfaces due to an interplay with Rashba spin-orbit coupling [76] (which we have neglected so far but discuss further below). Magnon-mediated pairing admits a qualitatively different depairing effect: Namely, $B_{\parallel}$ explicitly breaks the U(1) spin-rotation symmetry that the spin-canting phase breaks spontaneously—thereby generating a gap $\Delta_{\rm mag}\sim\sqrt{|B_{\parallel}|}$ for the Goldstone magnon mode. (Imagine the potential utility of an experimental knob for controllably gapping out phonons in conventional superconductors!) As explained in Appendix E.3, the field-induced gap regularizes the low-frequency divergence of the pairing interaction in Eq. (50) via the replacement $|\nu|\rightarrow|\nu|+\Delta_{\rm mag}$. The resulting dependence of $T_{c}$ on $B_{\parallel}$ poses a nontrivial problem worth future study, as it requires solving frequency-dependent Eliashberg equations, also accounting for the self-energy whose singularity at small ${\boldsymbol{q}}$ will be similarly cut off by $B_{\parallel}$. Nevertheless, we can at least qualitatively assert that our mechanism implies greater susceptibility to in-plane depairing effects compared to, say, a phonon-mediated scenario, possibly explaining peculiar Pauli-limit violation trends observed experimentally [37, 38, 40, 41].
@@
As alluded to above, Rashba spin-orbit coupling $\lambda_{R}$ generically accompanies Ising spin-orbit interaction in graphene multilayers proximitized by a TMD. And like in-plane magnetic fields, $\lambda_{R}$ explicitly breaks U(1) spin symmetry, gapping the Goldstone magnon even at $B_{\parallel}=0$. The bare Rashba energy scale is believed to be relatively weak, $\lambda_{R}\sim 1~{}\rm{meV}$, in graphene devices proximitized by WSe2 or WS2 [30, 31, 32, 33, 34, 35, 36, 73, 74, 75]. Crucially, however, the effective Rashba spin splitting in the low-energy bands strongly depends on the applied displacement field and the fermiology of the pockets relevant for pairing—and should be most prominent in large Fermi pockets. In our mechanism small Fermi pockets are favored (due to the $1/k_{F}$ dependence in the pairing interaction), which are predicted to be much less sensitive to Rashba depairing effects. In principle, measuring devices with variable $\lambda_{R}$, and ideally other parameters fixed, would provide another falsifiable check for our theory.
@@
We conclude with some speculation on extensions of our theory. Analyzing the favorability of spin-canting order in systems composed of more than three graphene layers, e.g., from a Hartree-Fock angle, would provide valuable insight into the wider applicability of our proposed pairing mechanism. It is also interesting to explore pairing mediated by Goldstone modes associated with other broken-symmetry orders [77], most notably intervalley coherence that also naturally emerges in graphene multilayers [12, 21, 47, 78, 50, 79, 51]. A particularly interesting possibility—which combines the two different types of continous symmetry breaking relevant for graphene systems—consists of intervalley coherent orders that also exhibit spin canting [51]. Such composite order can lead to generalized quarter-metal states that comprise a single, non-degenerate majority Fermi surface; similarly to the ideas developed in this work, our two-magnon exchange mechanism could potentially mediate Cooper pairing in this setting. Interestingly, one of the new superconducting regions in proximitized Bernal bilayer graphene [37] exhibits fermiology consistent with this description.
@@
Looking beyond graphene, can one identify a broader set of candidate materials for which Cooper pairing emerges from two-magnon exchange processes boosted by spin-orbit interaction? Ferromagnetic bulk materials composed of weakly coupled layers may be fruitful in this regard. In particular, the magnetic order could inherit a spin-canted structure upon introducing spin-orbit coupling; moreover, the layering could potentially allow magnons to mediate singular pairing interactions as in our analysis. For systems realizing more complex magnetic order (e.g., antiferromagnetism), can two-magnon exchange still provide a fruitful pairing mechanism—possibly leading to superconducting order parameters beyond simple $s$-wave? Finally, can one uncover excitonic analogues of our magnon-mediated pairing scenario?
@@ Acknowledgments
We are grateful to Trevor Arp, Nick Bultinck, Erez Berg, Andrey Chubukov, Long Ju, Hyunjin Kim, Patrick Lee, Leonid Levitov, Cyprian Lewandowski, Caitlin Patterson, Gal Shavit, Tomohiro Soejima, Alex Thomson, Yiran Zhang, Stevan Nadj-Perge, Owen Sheekey, and Andrea Young for insightful discussions. Z. D. and É. L.-H. are supported by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8682. Portions of this work were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (derivation of magnon spectrum and magnon-mediated interactions, J. A.). Additional support was provided by the Caltech Institute for Quantum Information and Matter, an NSF Physics Frontiers Center with support of the Gordon and Betty Moore Foundation through Grant GBMF1250.
@@ Appendix C Dispersion of magnons
with $b=s,a$ and $\eta=\pm$. We now solve for the low-energy magnon spectrum by examining the poles of these propagators. Specifically, we focus on the $s-$ and $a-$ branches since the former is gapless throughout the spin canted phase while the latter becomes gapless at the spin-canting phase transition. We generally ignore the gapped $s+$ and $a+$ branches (but see below for the special case of $\theta_{0}=\pi/2$). Solving the poles at small $q$ for the $s-$ mode, we find the following dispersion relation:
@@
Recalling the definition $\kappa=[4(V+J)\bar{m}^{2}]^{-1}$, we see that $c_{s}$ reduces to Eq. (29) from the main text in the physically motivated limit $J\ll V$. Similarly, for the $a-$ mode we obtain
@@
and is generally finite but decreases to zero as $\theta_{0}\rightarrow 0$.

### v3

@@ V Discussion
    The phase diagram area populated by superconductivity shrinks upon increasing spin-orbit coupling. In our scenario the pairing glue—soft magnons—relies on spin-canting order in the normal state. As $\lambda_{I}$ increases, however, canting order eventually loses in favor of an Ising spin-valley-locked phase; recall Sec. II.1. It is thus natural to anticipate that the spin-canted phase occupies a progressively smaller region of the phase diagram as $\lambda_{I}$ rises, as seen in Hartree-Fock studies [53, 54]. Our theory then predicts that superconductivity should also occupy a diminished area.
@@
    The critical temperature $T_{c}$ is enhanced upon increasing the spin-orbit coupling. This prediction follows directly from our analysis of the effective pairing strength in Eq. (45), which contains a $\cos^{2}\theta_{0}$ prefactor that is proportional to $\lambda_{I}^{2}$ through Eq. (4). It is then natural that turning on $\lambda_{I}$ increases the optimal $T_{c}$ emerging from a spin-canted normal state, even though the latter’s phase space diminishes.
@@
Remarkably, Ref. 42 observed both of these trends in bilayer devices, where the strength of Ising spin-orbit interaction was controlled through engineering the twist angle at the WSe2-graphene interface [63, 64, 65].
@@
Another nontrivial prediction concerns the dependence of superconductivity on Fermi surface structure. We showed that, due to its interband nature, our proposed pairing mechanism crucially relies on the presence of minority Fermi surfaces. More precisely, the dimensionless coupling constant in Eq. (46) depends on the minority carriers through (the square root of) their density of states. If modeling the minority pockets using a parabolic dispersion, we therefore expect superconductivity to be relatively insensitive to a decrease of minority carrier density, until it gets abruptly suppressed upon vacating the minority pockets. This prediction matches the generic behavior observed in bilayer and trilayer graphene, where prominent superconducting phases only appear in regions hosting minority pockets [40, 41, 42, 43, 44, 45]. Even more strikingly, in trilayers superconductivity abruptly terminates near Lifshitz transitions marking the disappearance of minority carriers [44, 45].
Direct detection of the spin-canting order central to our theory (e.g., through SQUID measurements, or by tracking the behavior of phase boundaries with applied magnetic fields in compressibility data) poses arguably the simplest falsifiable check for our proposed explanation. Perhaps the most striking consequence of pairing from a spin-canted normal state, due to its in-plane magnetic moment, is that the extent of superconductivity near phase boundaries to a non-magnetized competing phase can increase [^3] This behavior is not universal because $B_{\parallel}$ might also have competing pair-breaking effects, in particular as it breaks U(1) spin rotation symmetry and opens up a gap in the magnon dispersion, thus reducing the pairing interaction. Nevertheless, observing such an increase in phase space would constitute evidence for an in-plane magnetic moment in the superconducting phase. under the application of a small in-plane field. This behavior was recently reported in proximitized rhombohedral trilayers [44], along one of the phase boundaries that delimit the strongest spin-orbit-enabled superconductor.
@@
Our pairing scenario also has ramifications for depairing by in-plane magnetic fields $B_{\parallel}$. Conventional in-plane-field depairing mechanisms include orbital effects and warping of the Fermi surfaces due to an interplay with Rashba spin-orbit coupling [66] (which we have neglected so far but discuss further below). Since the pairing glue in our scenario is based on magnons, one might suspect a non-trivial magnetic-field dependence on $T_{c}$. Indeed, $B_{\parallel}$ significantly affects the magnon spectrum near ${\boldsymbol{q}}=0$ as it explicitly breaks the U(1) spin-rotation symmetry that the spin-canting phase breaks spontaneously—thereby generating a gap $\Delta_{\rm mag}$ for the Goldstone magnon mode. However, since our pairing mechanism utilizes magnons near $|{{\boldsymbol{q}}}|\sim k_{0}$ rather than ${{\boldsymbol{q}}}=0$, $B_{\parallel}$ leads to a weak suppression of the pairing glue strength, until the induced magnon gap $\Delta_{\rm mag}$ becomes comparable to the typical magnon energy $\Omega_{0}$ at $|{{\boldsymbol{q}}}|\sim k_{0}$. While our mechanism implies a greater depairing effect of in-plane magnetic fields compared to, say, a phonon-mediated scenario, its effect in a given platform will depend on whether orbital depairing, Rashba or magnon effects are predominant.
@@
As alluded to above, Rashba spin-orbit coupling $\lambda_{R}$ generically accompanies Ising spin-orbit interaction in graphene multilayers proximitized by a TMD. And like in-plane magnetic fields, $\lambda_{R}$ explicitly breaks U(1) spin symmetry, gapping the Goldstone magnon even at $B_{\parallel}=0$. The bare Rashba energy scale is believed to be relatively weak, $\lambda_{R}\sim 1~{}\rm{meV}$, in graphene devices proximitized by WSe2 or WS2 [33, 34, 35, 36, 37, 38, 39, 63, 64, 65]. Crucially, however, the effective Rashba spin splitting in the low-energy bands can be strongly suppressed by the applied displacement field and the fermiology of the pockets relevant for pairing. In principle, measuring devices with variable $\lambda_{R}$, and ideally other parameters fixed, would provide another falsifiable check for our theory.
@@
We conclude with some speculation on extensions of our theory. Analyzing the favorability of spin-canting order in systems composed of more than three graphene layers, e.g., from a Hartree-Fock angle, would provide valuable insight into the wider applicability of our proposed pairing mechanism. It is also interesting to explore pairing mediated by Goldstone modes associated with other broken-symmetry orders [67], most notably intervalley coherence that also naturally emerges in graphene multilayers [12, 21, 50, 68, 53, 69, 70, 54, 71, 72, 73, 74]. A particularly interesting possibility—which combines the two different types of continous symmetry breaking relevant for graphene systems—consists of intervalley coherent orders that also exhibit spin canting [54]. Such composite order can lead to generalized quarter-metal states that comprise a single, non-degenerate majority Fermi surface; similarly to the ideas developed in this work, our magnon-mediated pairing mechanism could potentially mediate Cooper pairing in this setting. Interestingly, one of the new superconducting regions in proximitized Bernal bilayer graphene [40] exhibits fermiology consistent with this description.
Another promising direction consists of investigating the spin-canting critical point ($\theta_{0}\rightarrow 0$) more closely. There, the two gapless ($a-$ and $s-$) modes may mediate pairing interactions with an interesting frequency dependence beyond BCS theory, which will be investigated in upcoming work [61]. This endeavor is further motivated by a recent experiment suggesting that superconductivity in spin-orbit-proximitized trilayers occurs on top of a phase transition where spin-canting order sets in [44].
@@
Looking beyond graphene, can one identify a broader set of candidate materials for which Cooper pairing emerges from magnon exchange processes boosted by spin-orbit interaction? Ferromagnetic bulk materials composed of weakly coupled layers may be fruitful in this regard. In particular, the magnetic order could inherit a spin-canted structure upon introducing spin-orbit coupling. For systems realizing more complex magnetic order (e.g., antiferromagnetism), can magnon exchange still provide a fruitful pairing mechanism—possibly leading to superconducting order parameters beyond a pseudospin-singlet $s$-wave? Finally, can one uncover excitonic analogues of our magnon-mediated pairing scenario?
@@ Acknowledgments
We are grateful to Trevor Arp, Erez Berg, Nick Bultinck, Andrey Chubukov, Julian Ingham, Long Ju, Hyunjin Kim, Patrick Lee, Leonid Levitov, Cyprian Lewandowski, Stevan Nadj-Perge, Caitlin Patterson, Gal Shavit, Owen Sheekey, Tomohiro Soejima, Alex Thomson, Yaar Vituri, Andrea Young, and Yiran Zhang for insightful discussions. We want to thank in particular Alex Thomson and Johnson Guanyao Chen for pointing out a mistake in the previous version of this manuscript, where an intraband two-magnon pairing interaction diverging as $q\rightarrow 0$ was obtained. This contribution is in fact canceled by higher-order terms in the magnon-fermion coupling; see footnote 2. Z. D. and É. L.-H. are supported by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8682. Portions of this work were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (derivation of magnon spectrum and magnon-mediated interactions, J. A.). Additional support was provided by the Caltech Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-2317110).
@@ Appendix C Dispersion of magnons
with $b=s,a$ and $\eta=\pm$. We now solve for the low-energy magnon spectrum by examining the poles of these propagators. Specifically, we focus on the $s-$ and $a-$ branches since the former is gapless throughout the spin canted phase while the latter becomes gapless at the spin-canted phase transition. We generally ignore the gapped $s+$ and $a+$ branches (but see below for the special case of $\theta_{0}=\pi/2$). Solving the poles at small $q$ for the $s-$ mode, we find the following dispersion relation:
@@
Similarly, for the $a-$ mode we obtain
@@
and is generally finite but decreases to zero as $\theta_{0}\rightarrow 0$. In comparison, both s+ and a+ mode have large gaps
$\Omega_{s+}=$ $\left[\frac{2\tilde{J}}{\kappa}\cos^{2}\theta_{0}+\frac{\gamma^{2}}{\kappa^{2}}\right]^{\frac{1}{2}}$ (62)
$\Omega_{a+}=$ $\left[\frac{\tilde{J}}{\kappa}(1+\sin^{2}\theta_{0})+\frac{\gamma^{2}}{2\kappa^{2}}+\sqrt{\left(\frac{\tilde{J}}{\kappa}(1+\sin^{2}\theta_{0})+\frac{\gamma^{2}}{2\kappa^{2}}\right)^{2}-\frac{4\tilde{J}^{2}}{\kappa^{2}}\sin^{2}\theta_{0}}\right]^{\frac{1}{2}},$ (63)
which are dominated by the term $\frac{\gamma^{2}}{2\kappa^{2}}=2\bar{m}^{2}$ and are thus comparable to the Fermi energy $\epsilon_{F}$. Therefore, these modes are expected to be irrelevant for pairing due to an absence of separation with the Coulomb energy scale, and a corresponding lack of retardation effects.